What is the solution for x= 3-y and -y=-5-x?
Answer by ewatrrr(24785)

 
Hi,
What is the solution for x= 3-y and -y=-5-x?
Putting each EQ given in terms of y:
y = 3-x
&y = x+5
3-x = x+5
-2 = 2x
-1 = x and y = 4